Input/Output Help

Suppose that in order to produce a good, Y, a firm needs three inputs, capital, K, labor, L, and finally, land, W. With these a firm can produce Y- K L w° where a+b+g 1. If a firm double all inputs, what will happen to his the output? Select one a. It will double. O b. More information is needed c. It will less than double. O d. It will more than double
Answer #1

Option (d).

Y = KaLbWg

If all inputs are doubled, new production becomes becomes

Y1 = (2K)a(2L)b(2W)g = 2a x 2b x 2g x KaLbWg = 2a x 2b x 2g x Y = 2(a+b+g) x Y

Y1/Y = (2(a+b+g) > 2, since (a + b + g) >1 [given].

[Since doubling all inputs more than doubles the output, there is increasing return to scale]